OpenStudy (anonymous):

Please help i don't understand what to do! write an equation in slope-intercept form of the line that passes through the given point and is parallel to the graph of the given equation. (1,-3) y+2=4(x-1) (0,0) y=2/3x+1

OpenStudy (anonymous):

@Catlover5925

OpenStudy (anonymous):

do you mind if i help you?

OpenStudy (anonymous):

no not all all thank you

OpenStudy (anonymous):

at*

OpenStudy (anonymous):

:) slope intercept form of a line is given by : \(\sf y=mx+b\), where \(\sf m\) is the slope and \(\sf b\) is the y-intercept. the given lines from your question are in the form \((y-y_1)=m(x-x_1)\), where \(\sf m\) is the slope So now, let's do the first one: \(\sf y+2=4(x-1) \) the question is asking for the line \(\sf parallel\) to the given line, which means that the required line will have the \(\sf same\ slope\) with the given line. so for the given line, what is the slope \(\sf m\)?

OpenStudy (catlover5925):